Database last updated on Tue, 27 Jan 2026 13:10:52 +0000 by localhost
Instrument constant changes for the week before 2025064 (3 sec, 0 rows)
SL burned out? (AP results when SL is ON) (0 sec, 11 rows)
Non-critical warnings from b-files' comments for days between 2025064 and 2025064 (0 sec, 1 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| rl: | 2025 064 05 43 44 Reset Warm -> Tepid from 2025 064 05 43 38 until 2025 064 05 43 39 . Counts: 1,0 | 1 | 3 | B06425.192 | TORON | #192 | 05:38:28 | | 1 | -1 |
Measurements and tests summary.
| _jday | _loc | _serial | _skc | dt | ed | fz | hg | hp | rs | sl | uv | zs |
| 064 | TORONTO | #192 | ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.097) | 2 | 2 | 8 | 11 | 11 | 2 | 8 | 24 | 77 |
(0 sec)
Relative and absolute humidity for day 2025064 (2 sec, 1 rows)
| Serial | RH, % | AH, g/m3 | 7-day RH history | Des-nt replcd days ago |
| 192 | 9.4 | 1.4 |